Yoo, Inc., has arranged a line of credit that allows it to borrow up to $40 million at any time. The interest rate is .616 percent per month. Additionally, the company must deposit 5 percent of the amount borrowed in a non-interest bearing account. The bank uses compound interest on its line-of-credit loans. If the company needs $16 million for 5 months, how much will it pay in interest? Multiple Choice

$623,636.01 $525,167.17 $583,519.08 $748,363.22 $498,908.81
Answer #1

Amount Required = $16 million

Amount Borrowed = 16/(1 - 0.05)

Amount Borrowed = $16,842,105.00

Amount to be paid back = 16,842,105(1.00616)5

Amount to be paid back = $17,367,272.17

Interest Paid = 17,367,272.17 - 16,842,105.00

Interest Paid = $525,167.17